OpenStudy (anonymous):

What is the value of n -1 of this data? 3, 12, 15, 25, 30, 40

OpenStudy (anonymous):

if n is the number of elements in the sample then n-1=6-1=5 are you doing t-distribution or chi-square?

OpenStudy (anonymous):

I have no idea

OpenStudy (amistre64):

they are learning how to recognize the properties required in statistics